Given that log4 (y-1) +log4(x/1)=1 and log2 (y+1)+log2 x=2,solve for x and y





1
Expert's answer
2022-01-31T16:06:49-0500

"(y-1)(x)=4" "......(i)"

"(y+1)(x)=2^2.....(ii)"


Dividing "(ii)" by "(i)"


"\frac{y-1}{y+1}=1"

"y-1=y+1"

"0=2"


"\therefore" No solution


Rewriting the question to read ;


log"_4(y-1)+" log"_4(\frac{1}{2}x)=1"

log"_2(y+1)+" log"_2x=2"


Then : Using laws of logarithms

"\frac{1}{2}x(y-1)=4....(i)"

"x(y+1)=4.....(ii)"


Dividing "(ii)" by "(i)"

"\frac{2(y+1)}{y-1}=1"


Solving

"y=-3"


Substituting in "(ii)"

"x(-3+1)=4"

"x=-2"